OpenStudy (anonymous):

please help on the algebra 2 problem!!

OpenStudy (anonymous):

And what is that?

OpenStudy (anonymous):

this is the question

OpenStudy (anonymous):

let's see... since the sides are the same... and it's a right triangle... \(2x^2=9^2\) u know how to get this part right?

OpenStudy (anonymous):

@sydneyarnea

OpenStudy (anonymous):

i have no clue @Molly_Qian

OpenStudy (anonymous):

ok u know the pathrgoran throm or whatever?

OpenStudy (anonymous):

the \(A^2+B^2=C^2\) thing?

OpenStudy (anonymous):

since bosh A and B are the same (x), it gives you something like \(X^2+X^2=9^2\)

OpenStudy (anonymous):

which gives you \(2x^2=81\)

OpenStudy (anonymous):

and now all you have to do is find x

OpenStudy (anonymous):

it depends on how you do it

OpenStudy (anonymous):

9 2?

OpenStudy (anonymous):

you can spuare root the whole thing and get you \(2x=9\) and then divide the whole thing by 2 which get you \(4.5\)

OpenStudy (anonymous):

which is the easiest way to do it

OpenStudy (anonymous):

so its a or b?

OpenStudy (anonymous):

apple